OpenStudy (anonymous):

Simplify the square root of x to the 4th power

OpenStudy (anonymous):

\[x^2\]

OpenStudy (anonymous):

\[x ^{2}\]

OpenStudy (anonymous):

\[\sqrt{x \sqrt[?]{4}}\]

OpenStudy (anonymous):

\[(\sqrt{x})^4=\sqrt{x}\sqrt{x}\sqrt{x}\sqrt{x}=xx=x^2\]

OpenStudy (anonymous):

nope...it is \[\sqrt{x ^{4}}=x^2\]

OpenStudy (anonymous):

nope intended at davette :)